# FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks

### v2

@@ 7 Conclusion
Given the significance of spatial reasoning in AI, we introduce Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark to evaluate Frame of Reference comprehension in textual spatial expressions through question-answering and grounding in visual modality via diffusion models. Using this benchmark, we identify notable differences in FoR comprehension across LLMs and their struggle with questions that require adaptation between multiple FoRs. Moreover, the bias in FoR interpretations impacts the layout generation with LLMs for text-to-image models. To improve FoR comprehension, we propose Spatial-Guided prompting, which extracts spatial relations’ topological, distal, and directional type information in addition to FoR and includes this information in downstream task prompting. Employing SG improves performance in both QA tasks requiring FoR understanding and text-to-image generation.
8 Limitations
While we analyze LLMs’ shortcomings, our benchmark only highlights areas for improvement. The trustworthiness and reliability of the LLMs are still a research challenge. Our analysis is confined to the spatial reasoning domain and does not account for biases related to gender or race. However, we acknowledge that linguistic and cultural variations in spatial expression are not considered, as our study focuses solely on English. Extending this work to multiple languages could reveal important differences in FoR adaptation. Our analysis is still limited to the synthetic environment. Future research should consider the broader implications of the frame of reference of spatial reasoning in real-world applications. Additionally, our experiments require substantial GPU resources, limiting the selection of LLMs and constraining the feasibility of testing larger models. The computational demands also pose accessibility challenges for researchers with limited resources. We find no ethical concerns in our methodology or results, as our study does not involve human subjects or sensitive data.
Appendix A Additional details of FoREST Dataset
@@
A.1 Dataset Statistics
The FoREST dataset statistic is provided in the Table 4 for A-split and Table 5 for C-split.
A.2 Details Creation of FoREST dataset
We define the nine categories of objects selected in our dataset as indicated below in Table 6. We select sets of locatum and relatum based on the properties of each class to cover four cases of frame of reference defined in Section 3.2. Notice that we also consider the appropriateness of the container; for example, the car should not contain the bus.
Based on the selected locatum and relatum. To create an A-split spatial expression, we substitute the actual locatum and relatum objects in the Spatial Relation template. After obtaining the A-split contexts, we create their counterparts using the perspective/topology clauses to make the counterparts in the C-split. Then, we obtain the I-A and I-C split by applying the directional template to the first occurrence of relatum when it has intrinsic directions. The directional templates are "that is facing towards," "that is facing backward," "that is facing to the left," and "that is facing to the right." All the templates are in the Table 7. We then construct the scene configuration from each modified spatial expression and send it to the simulator developed using Unity3D. Eventually, the simulator produces four visualization images for each scene configuration.
@@ A.2 Details Creation of FoREST dataset
Table 6: All selected objects with two properties: intrinsic direction, affordance of being container
A.2.1 Simulation Details
The simulation starts with randomly placing the relatum into the scene with the orientation based on the given scene configuration. We randomly select the orientation by given scene configuration, [-40, 40] for front, [40, 140] for left, [140, 220] for back, and [220, 320] for right. Then, we create the locatum from the relatum position and move it in the spatial relation provided. If the frame of reference is relative, we move the locatum based on the camera’s orientation. Otherwise, we move it from the relatum’s orientation. Then, we check the camera’s visibility of both objects. If one of them is not visible, we repeat the process of generating the relatum until the correct placement is achieved. After getting the proper placement, we randomly choose the background from 6 backgrounds. Eventually, we repeat the procedures four times for one configuration.
A.2.2 Object Models and Background
For the object models and background, we find it from the unity assert store [^2] https://assetstore.unity.com. All of them are free and available for download. All of the 3D models used are shown in Figure 4.
@@ A.2.2 Object Models and Background
A.2.3 Textual templates
All manually created templates used to create FoREST are given in Table 7.
In the camera view, how is {locatum} positioned in relation to {relatum}?
Based on the camera perspective, where is the {locatum} from the {relatum}’s position?
Question Templates From the camera perspective, what is the relation of the{locatum} to the {relatum}?
Looking through the camera perspective, how does {locatum} appear to be oriented relative to {relatum}’s position?
Based on the camera angle, where is {locatum} located with respect to {relatum}’s location?
Table 8: All templates used to create FoREST dataset.
A.2.4 GPT template
We prompt GPT with an initial question using a specific template (the first template in Table 8) to generate four additional variations. Each generated template was manually reviewed and validated before being added to the set of questions in our corpus. All templates are listed in Table 8. Question templates for asking the spatial relation from relatum view are edited by replacing references from the camera to the relatum object.
Appendix B Comparison between Qwen2 and Qwen2-VL in QA setting
@@
According to the A-split results in Table 1, Qwen2VL demonstrates a more balanced preference pattern and reasonable performance compared to Qwen2. This suggests that, unlike Qwen2, which often extracts spatial relations without accounting for perspective, Qwen2VL appears capable of reasoning about perspective shifts as well. This observation is further supported by the C-split results in Table 2, where Qwen2VL significantly outperforms Qwen2 in cases requiring adapting the spatial relations mentioned between different perspectives in most settings. These results support our hypothesis that visual training enhances the model’s ability to interpret perspective changes.
However, despite these improvements, Qwen2VL still struggles to provide reasoning from the relatum’s perspective on questions in CoT and SG+CoT settings. This suggests that while the model leverages visual training to reason about perspective and perform perspective shifts when necessary, its reasoning is more accurate when questions are framed from the camera’s viewpoint. However, it struggles with questions asked from the relatum’s perspective. This limitation is expected, as training data—particularly image captioning datasets—are typically annotated from a human, camera-based perspective, restricting the model’s ability to generalize across different viewpoints.
Appendix C Attention Analysis
@@
C.1 Experiment Setting
We conducted an attention-based interpretability analysis using Qwen2 in a 0-shot setting, where only the context and question were provided. We focused on analyzing attention weights across key semantic elements in both A-split and C-split. We analyzed the attention weights for various groups of tokens that convey the following concepts: locatum mention occurring in the context, relation in context, relatum in context, perspective in context, relatum orientation, perspective in question, locatum in question, and relatum occurring in the question.
We categorized response patterns into three scenarios: (1) the model answers incorrectly, (2) the model answers correctly with the assumption that the same perspective holds for context and question, and (3) the model answers correctly with the assumption that context and question have different perspectives. Note that we separate the analysis for both splits because the conceptual tokens are different, that is, the perspective mentioned in the context is only for the C-split.
C.2 Results
@@
Figure 7: Attention weights of input context for Question Answering in C-split. Considering the following semantic concepts: locatum mention in context, relation in context, relatum in context, perspective in context, relatum orientation, perspective in question, locatum in question, and relatum in question. Two patterns are considered: (1) incorrect answers, (2) correct answers. We separate the questions that require and not-require perspective changing.
@@
Appendix D Quantitative Examples
D.1 Example of QA Bias in the A-split
To illustrate example biases in the A-split, we present three types of model behavior. The first, shown in Figure 8, occurs when the model assumes a shared perspective between context and question, allowing it to extract spatial relations directly without FoR reasoning. This scenario is the most common behavior of Qwen2. The second, illustrated in Figure 9, arises when the model assumes the context always reflects the relatum’s perspective, favoring intrinsic over relative interpretation. Lastly, Figure 10 shows the model assuming the context always adopts the camera perspective.
Figure 8: The first example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model responds with the spatial relation in the context, assuming all questions and context have the same perspective.
Figure 9: The second example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model assumes the context has a relatum perspective.
Figure 10: The third example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model assumes the context has a camera perspective.
D.2 Incorrect QA reasoning in the C-split
Our results demonstrate that models struggle with converting between different perspectives when required to answer the questions, especially when the correct answer is either left or right. This difficulty contributes to their performance, as discussed in research question 2 in the paper (RQ2). To further illustrate LLMs’ struggle, we provide quantitative examples of failure cases, focusing specifically on the models’ most challenging instances in Figure 11.
Figure 11: An example of an incorrect SG+CoT case, where the model generates incorrect reasoning and mistakenly converts left to right.
Appendix E Template Variation in Question-Answering
@@
We report average accuracy on the C-split across all question templates for Qwen2 and GPT-4o in Table 9. While results show notable variation across templates, no single template consistently outperforms others across all tasks. Template effectiveness depends on both the prompting technique and the underlying model. Simpler templates generally yield higher accuracy, though more complex expressions may reduce overall performance. Nevertheless, the comparative trend remains consistent, as SG+CoT often outperforms CoT in overall improvement across both models. Notably, SG prompting yields the smallest performance variation across templates.
@@
Appendix F VISOR-uncond Score

### v3

@@ 7 Conclusion
Given the significance of spatial reasoning in AI models, we introduce the Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark to evaluate FoR comprehension in textual spatial expressions through question answering and grounding in the visual modality via diffusion models for text to image generation. Using this benchmark, we identify notable differences in FoR comprehension across LLMs, as well as their struggles with questions that require adapting between multiple FoRs. Moreover, biases in FoR interpretation affect layout generation in text-to-image models. To improve FoR comprehension, we propose Spatial-Guided prompting, which extracts topological, distal, and directional information in addition to FoR, and incorporates this knowledge into downstream prompting. Employing SG improves performance in both QA tasks requiring FoR understanding and text-to-image generation models by providing a more accurate layout to these models.
Acknowledgment
This project is partially supported by the Office of Naval Research (ONR) grant N00014-23-1-2417. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Office of Naval Research. We thank anonymous reviewers for their constructive feedback, which greatly helped us improve this manuscript.
Limitations
While we analyze LLMs’ shortcomings, our benchmark only highlights areas for improvement. The trustworthiness and reliability of the LLMs are still a research challenge. Our analysis is confined to the spatial reasoning domain and does not account for biases related to gender or race. We acknowledge that linguistic and cultural variations in spatial expression are not considered, as our study focuses solely on English. Extending this work to multiple languages could reveal important differences in FoR adaptation. Our analysis is still limited to the synthetic environment. Future research should consider the broader implications of the frame of reference of spatial reasoning in real-world applications. Additionally, our experiments require substantial GPU resources, limiting the selection of LLMs and constraining the feasibility of testing larger models. The computational demands also pose accessibility challenges for researchers with limited resources. We find no ethical concerns in our methodology or results, as our study does not involve human subjects or sensitive data.
Appendix A Additional Details of FoREST Creation
@@
We define the nine categories of objects selected in our dataset as indicated below in Table 5. We select sets of locatum and relatum based on the properties of each class to cover four cases of frame of reference defined in Section 3.2. Notice that we also consider the appropriateness of the container; for example, the car should not contain the bus. Based on the selected locatum and relatum. To create an A-split spatial expression, we substitute the actual locatum and relatum objects in the Spatial Relation template. After obtaining the A-split contexts, we create their counterparts using the perspective/topology clauses to make the counterparts in the C-split. Then, we obtain the I-A and I-C split by applying the directional template to the first occurrence of relatum when it has intrinsic directions. The directional templates are “that is facing towards,” “that is facing backward,” “that is facing to the left,” and "that is facing to the right.” All the templates are in the Table 6. We then construct the scene configuration from each modified spatial expression and send it to the simulator developed using Unity3D. Eventually, the simulator produces four visualization images for each scene configuration.
@@
Table 5: All selected objects with two properties: intrinsic direction, and affordance of being a container
A.1 Scene generation
The process begins by randomly placing the relatum in the scene with an orientation specified by the scene configuration. The relatum’s orientation is then sampled from predefined ranges: [-40, 40] for front, [40, 140] for left, [140, 220] for back, and [220, 320] for right. Next, the locatum is positioned relative to the relatum according to the given spatial relation. If the FoR is relative, the locatum is placed with respect to the camera’s orientation; otherwise, it is placed with respect to the relatum’s orientation. We then check whether both objects are visible from the camera. If not, the process regenerates the locatum and the relatum until a valid placement is achieved. Once placement is finalized, one of six backgrounds is randomly selected. This procedure is repeated four times for each scene configuration.
A.2 Object models and background
For the object models and background, we obtain them from the Unity Asset Store [^2] https://assetstore.unity.com. All of them are free and available for download. All 3D models used are shown in Figure 4.
@@ A.2 Object models and background
A.3 Templates
Context templates All manually created templates used to create a FoREST spatial expression are given in Table 6.
In the {perspective}, how is {locatum} positioned in relation to {relatum}?
Based on the {perspective}, where is the {locatum} from the {relatum}’s position?
Normal Templates From the {perspective}, what is the relation of the{locatum} to the {relatum}?
Looking through the {perspective}, how does {locatum} appear to be oriented relative to {relatum}’s position?
Based on the {perspective}, where is {locatum} located with respect to {relatum}’s location?
In relation to the {relatum}, where is the {locatum} located when viewed from {perspective}?
In the {perspective}, given {relatum} as reference, where is {locatum} located?
Reverse Templates Relative to the {relatum}, where can the {locatum} be found from the viewpoint of {perspective}?
Considering {relatum} as reference, where does {locatum} lie when seen from {perspective}?
Table 7: Question templates used to construct the FoREST dataset. Normal templates refer to cases where the locatum precedes the relatum, while reversed templates are the opposite. Normal templates are used in the main experiments, and reversed templates in the additional experiment (Appendix F.2).
Question Templates. To generate question templates, we first prompt GPT-4o with a manually created template (the first in Table 7). GPT-4o then generates eight additional variations, including versions with the relatum and locatum in reversed order. Each template was manually reviewed and validated before being added to our corpus; the full set is listed in Table 7. Templates from the camera’s perspective use the camera as the perspective, while templates from the relatum’s perspective use the relatum object. For each generated scenario, we select one normal and one reversed template to ensure both orders are represented, allowing us to evaluate whether the order of the locatum and relatum influences model performance. In the FoREST dataset, the contexts of both orders are kept separate.
Appendix B Impact of Visual Training
@@
To analyze the impact of visual training, we compare Qwen2 with Qwen2-VL, which extends Qwen2 with visual perception capability. In the A-split (Table 2), Qwen2-VL shows a more balanced preference pattern and stronger performance than Qwen2. This suggests that, unlike Qwen2, which often extracts spatial relations without accounting for perspective, Qwen2-VL can also reason about perspective shifts. This observation is further supported by the C-split results in Table 3, where Qwen2-VL substantially outperforms Qwen2 in cases requiring adaptation of spatial relations across perspectives. These findings reinforce our hypothesis that visual training improves a model’s ability to interpret perspective changes. Nevertheless, Qwen2-VL still struggles to reason from the relatum’s perspective in CoT and SG+CoT settings. While visual training enables it to handle perspective shifts more effectively, its reasoning remains more accurate when questions are framed from the camera’s viewpoint. This limitation is expected, since training data—particularly image-captioning datasets—are typically annotated from a human, camera-based perspective, restricting the model’s ability to generalize across viewpoints.
Appendix C GPT-o4-mini-high Results
For comparison with the main paper results, we include newer modelss such as GPT-4o-mini-high, a model configured for enhanced reasoning. As shown in Table 8, GPT-4o-mini-high exhibits a balanced preference between relative and intrinsic interpretations in the A-split. Similar to Qwen2VL, it demonstrates strong reasoning ability, achieving high accuracy even when FoR adaptation is required, likely due to its visual reasoning capability. Notably, in the C-split, GPT-4o-mini-high performs well on relatum-perspective questions, even outperforming GPT-4o (SG+CoT) in Table 9. However, its performance declines on camera-perspective questions, perhaps due to training that emphasizes alternative perspectives over the camera view. These results indicate that perspective shifting remains a challenge in the textual domain, even for strong reasoning models.
Question with Camera Perspective Question with Relatum perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
R% I% Acc. R% I% Acc. Acc. Acc. Acc. R% I% Acc. R% I% Acc. Acc.
Table 8: Additional results of QA accuracy in the A-Split with GPT-o4-mini-high. R% and I% represent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
Question with Camera Perspective Question with Relatum Perspective
GPT-4o (0-shot) $79.7$ $45.1$ $39.5$ $90.2$ $64.2$ $46.9$ $88.5$ $98.2$ $34.8$ $67.5$
Appendix D Attention Analysis
@@
D.1 Experimental setting
We conducted an attention-based interpretability analysis using Qwen2 in a 0-shot setting, where only the context and question were provided. We focused on analyzing attention weights across key semantic elements in both A-split and C-split. We analyzed the attention weights for various groups of tokens that convey the following concepts: locatum mention occurring in the context, relation in context, relatum in context, perspective in context, relatum orientation, perspective in question, locatum in question, and relatum occurring in the question. We categorized response patterns into three scenarios: (1) the model answers incorrectly, (2) the model answers correctly with the assumption that the same perspective holds for context and question, and (3) the model answers correctly with the assumption that context and question have different perspectives. Note that we separate the analysis for both splits because the conceptual tokens are different, that is, the perspective mentioned in the context is only for the C-split.
D.2 Experimental results
@@
Figure 7: Attention weights of input context for Question Answering in C-split. Considering the following semantic concepts: locatum mention in context, relation in context, relatum in context, perspective in context, relatum orientation, perspective in question, locatum in question, and relatum in question. Two patterns are considered: (1) incorrect answers, (2) correct answers. We separate the questions that require and do not require perspective changing.
@@
Appendix E Analysis of Various Spatial Configurations
In this section, we performed additional diagnostics by analyzing key features that may contribute to model failures, including the facing direction of the relatum of C-split and the type of spatial relations expressed in the context. This analysis is conducted over two settings: 0-shot and SG+CoT, using both Qwen2-72B and GPT-4o models.
E.1 Facing direction
Qwen2-72B + 0-shot GPT-4o + 0-shot
Facing Direction Front Back Left Right Avg. Facing Direction Front Back Left Right Avg.
Front 83.97 80.53 33.21 33.78 57.87 Front 84.54 77.29 66.79 69.85 74.62
Back 70.42 66.79 29.20 32.82 49.81 Back 58.02 47.52 16.41 30.15 38.02
Left 0.19 0.00 2.29 39.50 10.50 Left 96.95 61.64 5.73 9.92 43.56
Right 0.00 0.00 26.15 3.63 7.44 Right 64.31 26.34 1.53 1.72 23.47
Qwen2-72B + SG + CoT GPT-4o + SG + CoT
Facing Direction Front Back Left Right Avg. Facing Direction Front Back Left Right Avg.
Front 98.09 83.02 6.87 13.17 50.29 Front 50.95 58.78 95.80 95.99 75.38
Back 75.00 75.57 37.60 36.83 56.25 Back 86.26 93.89 2.86 2.29 46.33
Left 12.21 19.27 0.38 83.78 28.91 Left 73.47 35.11 51.34 63.17 55.77
Right 5.73 48.28 57.25 9.73 30.25 Right 79.01 56.49 60.31 49.81 61.40
Table 10: The accuracy of each relation type (column-label) for each facing direction (row-label) presented in the input context for Qwen2-72B and GPT-40 with 0-shot and SG+CoT settings.
According to Table 10, we observe that the facing direction plays a significant role in the performance drop for both Qwen2-72B and GPT-4o. Both models achieve relatively high performance only when the relatum is facing either toward or away from the camera, compared to when it is facing left or right relative to the camera. This effect is particularly noticeable in GPT-4o, which shows a substantial performance gap between cases where the relatum is facing the camera and other directions. We speculate that this error stems from biases in the training data, which typically consist of image–caption pairs based on 2D images where objects are clearly oriented toward the camera. As a result, LLMs appear to overfit to FoR reasoning patterns on specific facing directions (facing toward the camera), leading to confusion—particularly between “left” and “right” that usually reverse when the relatum is facing toward the camera—when the relatum is oriented differently, even if the image is not presented in the input.
E.2 Spatial relation
Qwen2-72B + 0-shot GPT-4o + 0-shot
Spatial Relation Front Back Left Right Avg. Facing Direction Front Back Left Right Avg.
Front 83.97 66.79 2.29 3.63 39.17 Front 84.54 47.52 5.73 1.72 34.88
Back 70.42 80.53 26.15 39.50 54.15 Back 58.02 77.29 1.53 9.92 36.69
Left 0.19 0.00 29.20 33.78 15.79 Left 96.95 26.34 16.41 69.85 52.39
Right 0.00 0.00 33.21 32.82 16.51 Right 64.34 61.64 66.79 30.15 55.73
Qwen2-72B + SG + CoT GPT-4o + SG + CoT
Spatial Relation Front Back Left Right Avg. Facing Direction Front Back Left Right Avg.
Front 98.09 75.57 0.38 9.73 45.94 Front 50.95 93.89 51.34 49.81 61.50
Back 75.00 83.02 57.25 83.78 74.76 Back 86.26 58.78 60.31 63.17 67.13
Left 12.21 48.28 37.60 13.17 27.81 Left 73.47 56.49 2.86 95.99 57.20
Right 5.73 19.27 6.87 36.83 17.18 Right 79.01 35.11 95.80 2.29 53.05
Table 11: The accuracy of each relation type (column-label) for each spatial relation (row-label) presented in the input context for Qwen2-72B and GPT-40 with 0-shot and SG+CoT settings.
According to Table 11, we observe that Qwen2-72B does not always extract the relation in C-split. However, it shows some evidence of extraction when the spatial lexical form is “front” or “back” as the accuracy is higher in these cases. In scenarios where the spatial lexical form is either “left” or “right,” the model appears to attempt reasoning (i.e., going beyond extracting the same relation mentions in the context), but often fails to produce correct results. For GPT-4o (multimodal), the model demonstrates overall better performance than Qwen2, except in cases that require conversion to “left” or “right.” This may be attributed to errors introduced by facing direction, where the model appears to rely too heavily on reasoning patterns conditioned on the object facing toward the camera. Nevertheless, these results also indicate that the SG+CoT approach provides clear improvements when reasoning is required, though confusion between “left” and “right” persists.
Appendix F Template Variations Analysis
F.1 Performance for each question template
@@
We examine the effect of each question template on the performance of the models. Table 12 reports average accuracy on the C-split across all question templates for Qwen2-72B and GPT-4o. While results show notable variation across templates, no single template consistently outperforms others across all tasks. Template effectiveness depends on both the prompting technique and the underlying model. Simpler templates generally yield higher accuracy, though more complex expressions may reduce overall performance. Nevertheless, the comparative trend remains consistent, as SG+CoT often outperforms CoT in overall improvement across both models. Notably, SG prompting yields the smallest performance variation across templates.
@@
F.2 Order variation in question templates
We include an additional template experiment to ensure that results are independent of the order of spatial entities (locatum and relatum) in the questions. A reversed template is shown in Table 7.
Results. According to Table 13 and Table 14, the average accuracy difference between the original benchmark templates and the reverse versions is relatively small across A-split, C-split (Table 13 and Table 14). In particular, the difference ranges from 1% to 4%, with an average of 2.76%. Detailed comparisons can be seen by examining the new C-split of Table 14 alongside Table 3 in the main paper. Specifically, for questions asked from the camera perspective, performance slightly declines when using reverse templates, with the largest drop observed in the 4-shot setting (a decrease of 3.6%). In contrast, questions asked from the relatum perspective show improvement with reverse templates, particularly in the CoT setting, where accuracy increases by 3.8%. Notably, the challenge of perspective conversion persists and even worsens with reverse questions, suggesting that these gains are primarily due to cases where the model extracts spatial relations directly from the context without considering perspective. Finally, we observe that Qwen2 exhibits consistent behavior across both question orders by comparing the new results of A-split with Table 1. Qwen2 still favors answering FoR questions with the spatial lexicon explicitly.
Question with Camera Perspective Question with Relatum Perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
R% I% Acc. R% I% Acc. Acc. Acc. Acc. R% I% Acc. R% I% Acc. Acc.
Qwen2-72B (1) $70.6$ $29.0$ $94.1$ $69.0$ $30.6$ $92.6$ $100.0$ $100.0$ $94.7$ $26.5$ $73.1$ $85.5$ $27.1$ $72.5$ $87.6$ $85.8$
Qwen2-72B (2) $68.6$ $31.0$ $87.7$ $65.9$ $33.7$ $85.1$ $100.0$ $100.0$ $89.1$ $28.7$ $70.8$ $81.7$ $24.6$ $74.9$ $86.3$ $82.3$
Qwen2-72B (3) $45.3$ $54.4$ $86.1$ $41.6$ $58.2$ $85.5$ $100.0$ $100.0$ $88.0$ $35.1$ $64.5$ $87.8$ $36.6$ $63.0$ $85.7$ $87.5$
Qwen2-72B (4) $54.5$ $45.2$ $88.9$ $52.4$ $47.3$ $90.6$ $100.0$ $100.0$ $90.7$ $42.3$ $57.4$ $78.3$ $35.6$ $64.0$ $82.8$ $78.9$
Table 13: QA accuracy in the A-split with Qwen2-72B using templates where the locatum and relatum order is reverse, across all settings. R% and I% represent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
Question with Camera Perspective Question with Relatum Perspective
Qwen2-72B (0-shot) $90.6$ $32.9$ $31.8$ $91.1$ $64.0$ $26.4$ $96.8$ $99.8$ $24.4$ $61.7$
Qwen2-72B (4-shot) $87.3$ $33.9$ $34.4$ $72.6$ $61.7$ $31.0$ $96.7$ $99.4$ $22.7$ $63.5$
Qwen2-72B (CoT) $80.1$ $54.2$ $55.7$ $66.9$ $67.4$ $35.7$ $96.0$ $97.5$ $27.0$ $65.4$
Qwen2-72B (SG+CoT) $96.6$ $41.7$ $31.2$ $93.7$ $70.6$ $44.0$ $92.1$ $94.7$ $35.0$ $67.7$
Table 14: QA accuracy in the C-split with Qwen2-72B using templates where the locatum and relatum order is reverse, across all settings. ER, EI, II, and IR denote external relative, external intrinsic, internal intrinsic, and internal relative FoRs. Avg represents the micro-average accuracy. CP refers to context with camera perspective, while RP denotes context with relatum perspective.
Appendix G Human bias in Ambiguous Cases
Instruction:
G.1 Experimental setting
We provided a Google Form with instructions (Listing 1), followed by 150 questions. These were sampled from relatum categories in the A-split: 25 per case, covering four cases (Cow, Car, Box, Pen) for camera-perspective questions, and two cases (Cow, Car) for relatum-perspective questions, as Box and Pen lack intrinsic direction. No time constraints were imposed, and all participants were compensated at the standard research assistant rate. At the end, participants were informed about the study details and asked for consent to use their responses in the analysis. Results were evaluated using the metric defined in Section 5.1, and participant accuracies were averaged to establish the human baseline.
G.2 Experimental results
Individual human results are shown in Table 15. These results indicate that humans rely heavily on their assumptions when interpreting ambiguous frame-of-reference (FoR) contexts. Most participants favored the intrinsic FoR in A-split scenarios, though this preference varied across individuals, suggesting that background or prior visual experience may influence FoR interpretation. This aligns with findings from cognitive studies on how humans describe spatial relations, which indicate that recent examples may influence Coventry et al. (2018). However, our results contrast with other studies suggesting that humans tend to adopt an egocentric (relative) FoR in scenes with fewer landmarks and an allocentric (intrinsic) FoR in those with more landmarks Forster et al. (2023). Since our study does not provide visual input, we cannot infer how participants mentally reconstruct scenes to answer the questions with only two objects in the scene. The next notable point is that while Box and Pen cases appear straightforward for models (Table 2), humans may still struggle with questions in these categories. Humans do not exhibit the same limitations as LLMs in perspective-shifting tasks. When the context and question differ in perspective, humans still achieve over 90% accuracy, whereas LLMs often struggle with this type of reasoning.
Question with Camera Perspective Question with Relatum Perspective
Model Cow Car Box Pen Avg. Cow Car Avg.
R% I% Acc. R% I% Acc. Acc. Acc. Acc. R% I% Acc. R% I% Acc. Acc.
Table 15: QA accuracy in the A-Split of human study. R% and I% represent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
Appendix H Qualitative Observations
Example of QA Bias in the A-split. To illustrate example biases in the A-split, we present three types of model behavior. The first, shown in Figure 8, occurs when the model assumes a shared perspective between context and question, allowing it to extract spatial relations directly without FoR reasoning. This scenario is the most common behavior of Qwen2. The second, illustrated in Figure 9, arises when the model assumes the context always reflects the relatum’s perspective, favoring intrinsic over relative interpretation. Lastly, Figure 10 shows the model assuming the context always adopts the camera perspective.
Figure 8: The example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model responds with the spatial relation in the context, assuming all questions and context have the same perspective.
Figure 9: The example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model assumes the context has a relatum perspective.
Figure 10: The example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model assumes the context has a camera perspective.
Incorrect Reasoning in the C-split QA Task. To demonstrate LLMs’ confusion in left–right scenarios, we present a quantitative example of a failure case in Figure 11.
Figure 11: An example of an incorrect SG+CoT case, where the model generates incorrect reasoning and incorrectly converts left to right.
Appendix I VISOR-uncond Results